(-3,_____) is a solution of y=4-3x. Can you show me the step by step to solving this problem? Thank you
Answer by Edwin McCravy(20081)

The first coordinate of (-3,___) is given. The first coordinate
is what you substitute for x in order to get the value of y,
which is the second coordinate. So instead of x in the equation
y = 4 - 3x
we write (-3) in place of x, like this:
y = 4 - 3(-3)
and then simplify
y = 4 + 9
y = 13
And that value of y is the second coordinate and so the
answer is:
(-3,13)
Edwin